OpenStudy (anonymous):

What are the foci of the Ellipse? 36x^2+9y^2=324

OpenStudy (anonymous):

You can figure this out by writing the equation in ellipse form.

OpenStudy (anonymous):

So you have to divide both sides by 324.

OpenStudy (anonymous):

why 324? Because 324 divides evenly by both 9 and 36

OpenStudy (anonymous):

afterwards your equation should look like this \[x^2/9 + y^2/36 =1\]

OpenStudy (anonymous):

\[c^2=a^2-b^2\]

OpenStudy (anonymous):

But what are a and b?

OpenStudy (anonymous):

a^2 is the number under y^2. b^2 is the number under x^2

OpenStudy (anonymous):

So c=sqrt(36-9)

OpenStudy (anonymous):

c=5. Because c=5, the foci are located at (0,5) and (0,-5)

OpenStudy (anonymous):

http://www.mathwarehouse.com/ellipse/images/focus_equation_better.gif heres a picture if you still need help

OpenStudy (anonymous):

Okay Thankyou so much helped a lot!